Youre considering investing $1000 into a savings account at one of two possible banks. Youll leave the money invested for 2 years. The first bank (Bank A) is advertising a rate of 3% APR with monthly compounding. Unfortunately, Bank B hasnt washed their windows lately, and all you can see is that they compound interest daily (You cant see the APR or EAR rates for Bank B). Determine the APR associated with Bank B so that you would be indifferent between investing in the two banks.
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins s Cut aCopy Σ AutoSum ー E ゴWrap Text B า 프 . Ej-., Δ. : r_一 逻锂函Merge & Center. $, % , 弼,8 Paste Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard SM20 SK Alignment Number Cells Edting SL SM SN SO SP SQ SR ST SU SV SW SX SY BANK A RATE 3% MONTHLY COMPOUNDING 4 5 6 EAR- (1+r/m)^m-1- (1+0.03/12)^12-1-3.04160% BANK B RATE r96 DAILY COMPOUNDING 9 WE HAVE TO CALCULATE RATE SUCH THAT BOTH HAVE EQUAL EFFECTIVE RATE 10 EAR = 0.030416(1+/365)A365-1 1.030416(1+r/365)A365 1.03041641/365) = (1+r/365) 1.000082093 (1+r/365) 0.000082093 r/36!5 12 13 14 15 16 17 18 19 20 21 /4 | CALCULATOR .. LOAN OPTIONS r0.029964 [0.000082093 365] r 2.9964% CAP STRU VALUE BOX DILUTIONDOLLAR COST AVGMORTGAGE EXPO SheetiSheet2Shel 福 130% rences: x261 13:20 28-01-2019